OpenStudy (anonymous):

use the difference of squares to factor the expression 36x^8-9y^4

OpenStudy (anonymous):

36

OpenStudy (turingtest):

yeah, that's nonsense above

OpenStudy (anonymous):

Hi Turing! I understand factoring to a point, but not this specific method, would you be willing to walk me through it? :)

OpenStudy (anonymous):

wait, this should be in the math section right?

OpenStudy (anonymous):

oh I actually thought I put it in the math section! Completely my bad!

OpenStudy (turingtest):

lol I just followed the link, I didn't even know where we were :P

OpenStudy (turingtest):

difference of squares is the formula\[a^2-b^2=(a-b)(a+b)\]so you need to identify: what is a and b in you problem compare and contrast: the form\[a^2-b^2=(a+b)(a-b)\]and yours\[36x^8-9y^4\]what are \(a\) and \(b\) in your particular case?

OpenStudy (anonymous):

Sorry guys! I was talking in math chat and didn't pay attention to the actual page I was on!

OpenStudy (turingtest):

honest mistake, no worries :)

OpenStudy (anonymous):

a =36 and b= 9 ?

OpenStudy (turingtest):

good effort, but notice a few things...

OpenStudy (turingtest):

we need to keep the variables, so don't just drop x and y also not that I asked you what \(a\) and \(b\) are, but the formula starts with \(a^2\) and \(b^2\) so in your case, \(a^2=36x^8\) so what is \(b^2\) ?

OpenStudy (turingtest):

also note*

OpenStudy (anonymous):

b^2 = 9y^4 ?

OpenStudy (turingtest):

yes :) so if a^2=36x^8 and b^2=9y^4 then what are a and b ?

OpenStudy (anonymous):

I dont know :( would this be where I start factoring?

OpenStudy (anonymous):

would I take the square root of 36x^8? im so bad at this. but im thinking haha.

OpenStudy (turingtest):

yes, you just take the square root if\[a^2=36x^8\]then \[a=\sqrt{36x^8}\]

OpenStudy (turingtest):

simplify...

OpenStudy (anonymous):

6 sq root(x^8) sorry I dont know how to make the square root symbol

OpenStudy (turingtest):

that's fine, but what is \[\sqrt{x^8}\] ??? you can simplify that

OpenStudy (anonymous):

x^4?

OpenStudy (anonymous):

6x^4 so then b = 3y^2?

OpenStudy (turingtest):

you got it :) now just plug those values of a and b into the formula for difference of squares I wrote above

OpenStudy (anonymous):

36x^8 - 9y^4 = (6x^4 -3y^2) (6x^4 +3y^2) ? :)

OpenStudy (turingtest):

yep :) good job! think you got the idea ?

OpenStudy (anonymous):

yes! thanks so much! im actually super excited that I get it now!

OpenStudy (turingtest):

that's all I can hope for :)

OpenStudy (turingtest):

welcome!